Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Test-Stand  Not logged in ELOG logo
Message ID: 1047     Entry time: Tuesday, March 11, 2025, 14:36
Author: Brady Borrmann 
Target/No.#:  
Source:  
Station:  
Module:  
Target Oven W.O#: 44691-8 
Heat Shield W.O#: 44378-2 
Status: In progress 
Subject: Mass Marker Test Target 

Tuesday, March 11, 2025, 14:36 Roughing. Target installed.

Tuesday, March 11, 2025, 15:06 TP1B turned on.

Tuesday, March 11, 2025, 15:10 TP1A turned on. 

Tuesday, March 11, 2025, 15:18 IG1 Turned on 6.25e-05T 

Tuesday, March 11, 2025, 16:39 Turned water on. Started autoramping TGHT/TBHT to 580/230A at 1A/30s

Wednesday, March 12, 2025, 11:03  slowly ramping BIAS/EE/EZL to 30/2.5/13 kV 

Wednesday, March 12, 2025, 14:00 no beam seen out. It seems that FC0 is working though. Investigation needed

Wednesday, March 12, 2025, 16:00 decreased HV to 0 to investigate. See report (to be released and attached) for more information

Friday, March 14, 2025, 09:20 :  increasing TGHT TBHT to 100A at 1A/30s, then to 580/230A at 1.7A/10s. 

Friday, March 14, 2025, 11:31 increasing BIAS/EE/EZL to 30/2.5/13 kV 

 Friday, March 14, 2025, 12:57    it appears that everything is stable. FC0 = 3.6e-10 A (VQSX).  beginning mass scan.  mass scan saved as '2025-03-14-13-30 MM a 0A.txt'

Friday, March 14, 2025, 14:12 mass separator focused at mass 115 to visualize indium.

Friday, March 14, 2025, 14:12  increasing MM current, following mass peak.

__ Current __ FC6 output __

    0A        1.9e-11

    1A        2e-11

    2A        2.22-11

    3A        4.6e-11

    4A        5.9e-10

    5A        3.6e-9

    6A        8.6e-9

 Friday, March 14, 2025, 14:39  conducting mass scan.   '2025-03-14-15-20 MM b 6A.txt'   after the scan, it was noticed that the mm current had failed at some point during the scan.

Friday, March 14, 2025, 15:27   conducting mass scan with 0A mm current.  saved as 2025-03-14-15-55 MM c 0A.txt

Friday, March 14, 2025, 15:55 ramping down HV

Friday, March 14, 2025, 15:56 ramping down TGHT TBHT to 0 (1A/30s) to assess and correct mm OL issue

Tuesday, March 18, 2025, 14:49 After bringing down vacuum to open test stand, the mass marker connection has been investigated. Two multipin connections have become disconnected, seems to be at a soldered junction.

Tuesday, March 18, 2025, 14:51    MM has been moved to different pins for operation. multipin will be fixed at a later date. Now test stand has been vacuumed down and TGHT TBHT brought to 580 230 for measurement.

Tuesday, March 18, 2025, 14:51 MM temperatures:

MM current    _____    temperature:

0A ___ 195.8°C

1A ___ 198.1

2A ___ 223.7°C

3A ___ 282.1°C

4A ___ 370.1­­­°C

5A ___ 486.4°C

6A ___ 631.6°C

 

Tuesday, March 18, 2025, 16:17 cooling TGHT TBHT at -1A/30s to 0A

Wednesday, March 19, 2025, 16:30 :  Mass marker replaced - new ceramic oven and sleeve, new reservoir with Indium - 0.102g.  Same transfer tube and target assembly

Wednesday, March 19, 2025, 16:45:  Install complete, roughing vacuum

Wednesday, March 19, 2025, 17:07  Turbopumps on

Thursday, March 20, 2025, 10:00 :  Water cooling on. Increasing TGHT TBHT to 580 230 at 1A/30s.

Thursday, March 20, 2025, 12:00 :  at ~200A 200A, increased heating rate to 1A/10s

Thursday, March 20, 2025, 13:45 increasing HV BIAS/EE/EZL to 30/2.5/13 kV 

Thursday, March 20, 2025, 14:32   at HV now, stable. Beginning mass scan.    

 Thursday, March 20, 2025, 15:15  magnet positioned at mass 115

__ Current __ FC6 output __

    0A        1.17e-11

    1A        1.23e-11

    2A        2.60e-11

    3A        9.2e-11

    4A        2.8e-10

    5A        5.5e-10

    6A        1.4e-9

Thursday, March 20, 2025, 15:57  running mass scan - 2025-03-20_16-26 mm e 6A

 Thursday, March 20, 2025, 17:00 run mass scan - 2025-03-20_16-58 mm f 0A.txt
 
Thursday, March 20, 2025, 17:29  mm has cooled more, repeating mass scan.  2025-03-20_17-51 mm g 0A.txt
 
Thursday, March 20, 2025, 17:58  BIAS/EE/EZL down to 0V .  left hot for mm temperature evaluation tomorrow

Friday, March 21, 2025, 11:08   Temperature testing

Current ___ Temperature

0A     184.5

1A     187.2

2A     208.3

3A     268.7

4A     374.1

5A     419.2

6A     495.8

 Friday, March 21, 2025, 12:09   cooling TGHT TBHT at 1.7A/10s to make Faraday cage accessible for controls work

                           decreased cooling rate to 1A/20s for final 100A. at 14:00, test stand was opened for controls work.

Wednesday, March 26, 2025, 10:08   Turbo pumps switched off for venting and changing the mass marker reservoir

Wednesday, March 26, 2025, 18:01  Chamber opened, MMR replaced with MMR3, roughed, now turbopumps on.

Thursday, March 27, 2025, 12:38 TGHT TBHT brought up to 580 230 at 1A/30s for 100A and then 1.7A/10s for the rest. now increasing  HV BIAS/EE/EZL to 30/2.5/13 kV 

 Thursday, March 27, 2025, 13:58    at high voltage, beginning mass scan.   (h)

Thursday, March 27, 2025, 14:37    Magnet at 115, measuring beam output with increases of MM current

current ___ FC6 output

0A    8.2e-11

1A    8.2e-11

2A    8.6e-11

3A    1e-10

4A    1.88e-10

5A    5e-10

6A    8.7e-10

Thursday, March 27, 2025, 15:26 beginning mass scan at MM 6A. --- MM Current failed at the beginning of the mass scan. data still collected - indium peak lower than expected due to gradual cooling. mass scan saved (i) 2025-03-27_15-32 mm i 6A-0A.txt

Thursday, March 27, 2025, 15:33   cooled 0A mass scan beginning - (j) 2025-03-27_16-13 mm j 0A.txt

Thursday, March 27, 2025, 16:32  decreasing HV to 0

Friday, March 28, 2025, 09:13  spinning down turbo pumps to vent system. Must *begin* to repair/replace the multipin connection. - will affix twisted pair to two appropriate channels for IGLIS AC

Tuesday, April 08, 2025, 12:26   spinning down turbo pumps 

 Wednesday, April 09, 2025, 15:36  Twisted pair wire shielded, routed and crimped for iglis and high(er) current purposes. multipin channels 8 and 9.

                 roughing.

 Wednesday, April 09, 2025, 16:26   turbo pumps on

Thursday, April 10, 2025, 09:43    beginning heating - TGHT TBHT to 580 230 at 1A/30s

Thursday, April 10, 2025, 10:41     at 100A100A. increasing ramp rate to 1.7A/30s

Thursday, April 10, 2025, 11:39 increasing HV BIAS/EE/EZL to 30/2.5/13 kV 

Thursday, April 10, 2025, 13:12   -  mass scan - 0A mm. - -  2025-04-10_13-46 mm k 0A.txt
Thursday, April 10, 2025, 13:50     magnet on 115 u - turning on MMR current.

current ___ FC6 output

0A    1.56e-11

1A    1.5e-11

2A    1.81e-11

3A    2.1e-11

4A    3.1e-11

5A    5.2e-11

6A    1.75e-10

7A    1.08e-9

8A    5.3e-9

 

 Thursday, April 10, 2025, 14:49   mass scan at 8A. Frequent sparking, ~1/min but doesn't appear to be affecting anything. no affect on optics or power supplies.  appears to be inside TS chamber as well.   mass scan saved (l 8A) 2025-04-10_15-04 mm l 8A.txt

Thursday, April 10, 2025, 14:50   MMR current decreased to 6A. waiting to cool to begin mass scan. sparks have stopped .

Thursday, April 10, 2025, 15:08  beginning mass scan.  (m 6A)  saved. 2025-04-10_15-34 mm m 8A.txt

Thursday, April 10, 2025, 15:34  MMR current 0A. repeat mass scan after cooled. (n 0A) saved.  2025-04-10_16-10 mm n 0A.txt

Thursday, April 10, 2025, 16:02 voltages to 0 for temperature testing

Current ___ Temperature

0A    205     

1A     205.5 

2A     227    

3A     268    

4A     327    

5A     391    

6A     462      

7A     546

8A      645.6

Thursday, April 10, 2025, 16:51    cooling TGHT TBHT at 1A/30s

 

 Friday, April 11, 2025, 09:00  spinning down turbo pumps to vent chamber. must replace source cup (p2n style --> old style)

Friday, April 11, 2025, 09:56  opened VV1 to begin venting the chamber. 

Friday, April 11, 2025, 15:39    P2N source cup removed, old source cup installed for repeat test. roughing chamber now.

 Friday, April 11, 2025, 16:05   turbo pumps on

Monday, April 14, 2025, 09:43  Water cooling on. TGHT TBHT ramping up at 1A/30s to 580A/230A

Monday, April 14, 2025, 12:57 increasing HV BIAS/EE/EZL to 30/2.5/13 kV 

 Monday, April 14, 2025, 15:27   FC0 signal very low. running mass scan - (o - 0A MMR),  2025-04-14_15-42 mm o 0A.txt
Monday, April 14, 2025, 15:45   magnet set to 115u
increasing mmr current.

current ___ FC6 output

0A     2.4e-12

1A     4e-12     

2A     3.9e-12  

3A     4.8e-12

4A     3e-12 

5A    4e-12  

6A    4e-12

7A    4e-12

8A   4e-12

Monday, April 14, 2025, 16:06     no indium out was seen here due to increase of mmr current.

running mass scan again at 8A

MMR power supply tripped after 20u, had to reset power supply trip, quickly ramped down HV, reset, ramped back up.

mass scan running at 8A MMR.   appears that no beam is coming from this target...    spark occured halfway through the run that tripped the MMR power supply as well as the beam line vac gauge IG6 -> leads to interlock trip on many optics. Restored as quickly as possible. however, beam was far past 115, where indium had been expected    saved as  2025-04-14_16-42 mm p 8A.txt
 

 Monday, April 14, 2025, 16:44   dropping HV to 0V and cooling TGHT TBHT to 0A/0A for target examination

Monday, April 14, 2025, 16:53   This trip has caused some sort of communication issue with the TGHT, This caused TGHT to be 'stuck' at 580A. I had to eventually cut power to the supply and have power cycled it. have not been able to restore power to it remotely

Fixed TGHT. Was able reset it locally after opening the cage.

Tuesday, April 15, 2025, 08:06   Turbo pumps off to vent, remove target and examine. 

Tuesday, April 15, 2025, 08:43   VV1 open 

Tuesday, April 15, 2025, 11:22   Target opened, taken apart, examined. Appears that the MMR transfer line had become dislodged from the target block, preventing transfer. Fixed this, reassembled, put back in TS.

Tuesday, April 15, 2025, 11:23  roughing.

Tuesday, April 15, 2025, 12:00  turbo pumps on.

Tuesday, April 15, 2025, 13:23  water cooling on. no effect on vacuum

Wednesday, April 16, 2025, 09:37 ramping TGHT TBHT 

ELOG V2.9.2-2455